# PROPS: Progressively Private Self-alignment of Large Language Models

### v1

@@ 1 Introduction
The process of aligning LLMs relies on datasets comprising prompts, LLM-generated responses, and preference labels that indicate which response aligns better with human values, collectively referred to as preference data. The alignment approaches, including reinforcement learning with human feedback (RLHF) (Stiennon et al., 2020) and direct preference optimization (DPO) (Rafailov et al., 2024), leverage these preference datasets with ranked labels provided by human annotators.
@@
Motivation for Human Preference Privacy: Relying on human feedback for alignment brings up privacy concerns, particularly around the extent to which a labeler’s preferences might unintentionally expose their underlying values, identity, or personality. In medical applications, for example, publicly available case reports or anonymized symptoms can be used to prompt LLMs to generate diagnoses, with human feedback aligning these outputs to clinical best practices (Reddy, 2023; Han et al., 2024). Here, the preference data represents expert judgment about nuanced medical decisions and must remain private to protect the integrity and confidentiality of clinical expertise. In Wang et al. (2024), a large number of consultation records were used to construct a dataset to fine-tune an LLM to act as a chatbot physician, with GPT-4 used to retain records that highlighted professionalism, explainability, and emotional support. As the model was shown to be effective, this indicates that certain preferences of a doctor’s decision-making can be identified which could lead to potential exposure of a doctor’s preferences. Also, in policy analysis (Rao et al., 2023), publicly available survey questions or proposals can be used to elicit LLM-generated analyses, where policymakers’ feedback reveals sensitive interpretative insights that may require protection. Bakker et al. (2021), for example, observed that placing more emphasis on politics in their surveys to participants “resulted in self-reports of personality traits that were in some cases more aligned with preexisting political preferences.”
@@
Main Contributions: Motivated by the above observations, we study the problem of aligning LLMs with preference privacy. Specifically, we investigate two notions of privacy: a) preference-level privacy, and b) labeler-level privacy. Preference level privacy ensures that for any tuple $(x,y_{1},y_{2})$, where $x$ denotes the prompt and $y_{1},y_{2}$ denote the LLM-generated responses, the individual human-preference $\ell^{*}$ (which denotes whether $y_{1}$ or $y_{2}$ is preferable) does not significantly impact the aligned model. Formally, we leverage the existing notion of Differential Privacy (DP) Dwork et al. (2014), and use it to formalize the notion of $(\epsilon,\delta)$-preference-level DP, where $(\epsilon,\delta)$ represent the privacy budgets. The notion of labeler-level privacy (also commonly referred to as user-level privacy in the DP literature) hides the presence/absence of any individual human labeler and protects all the labels annotated by the labeler. We summarize and highlight the main contributions and novel aspects of this work:
  • PROPS for Private Alignment. We introduce Progressively Private Self-Alignment (PROPS), a multi-stage algorithm for alignment that improves privacy and utility. Instead of processing the entire perturbed dataset at once, PROPS divides alignment into stages. In the $k^{\text{th}}$ stage, for $(k>1)$, the model from the previous stage ($M_{k-1}$) is used along with non-private prompts, responses from a new batch, and noisy labels perturbed using Randomized Response (RR) $\ell_{\text{RR}}$. $M_{k-1}$ generates its own rankings $\ell_{M_{k-1}}$ which, combined with $\ell_{\text{RR}}$, are used to compute maximum-likelihood estimates (MLEs) for alignment. This staged process leverages intermediate models to improve preference labeling and reduce reliance on noisy labels, enhancing alignment quality while maintaining privacy. As the model evolves over $K$ stages, alignment becomes progressively refined. PROPS effectively balances privacy preservation with performance, offering a novel alignment framework.
  • Theoretical Insights. We study the utility-privacy tradeoffs for PROPS by analyzing the Sub-optimality gap Chowdhury et al. (2024a) defined as the difference between between the weights of a non-privately trained model with the weights of a privately trained model with a privacy budget $\epsilon$ in Section 3.1 (Theorem 1). The upper bound on this gap shows that PROPS (which uses MLE combining in the second stage) is no worse than vanilla RR and performs better as long as the intermediate model gets better at predicting preferences.
@@
LLM Alignment: Training LLMs typically involve three key stages: pre-training, supervised fine-tuning, and alignment. Among these, alignment is particularly important as it guides LLMs to produce responses that align with societal norms and human preferences. The alignment process relies on a dataset $\mathcal{D}$ consisting of $n$ samples, each containing a prompt $x$, LLM-generated responses $(y_{1},y_{2})$, and a human-preferred label $\ell^{*}$, collectively referred to as preference data. Two conventional methods for alignment, Reinforcement Learning with Human Feedback (RLHF) (Stiennon et al., 2020) and Direct Preference Optimization (DPO) (Rafailov et al., 2024), utilize preference datasets with ranked labels provided by human annotators. While these aligned models improve the quality of generated responses, they introduce privacy risks, particularly concerning the identity of the annotators and their associated preferences. Recent work shows membership inference attack on preference data for LLM alignment (Feng et al., 2024), and highlights the vulnerabilities of using human annotated preference data during alignment. To mitigate the privacy risks associated with human-annotated preference data, the notion of differential privacy (DP) (Dwork et al., 2014) has recently been explored for fine-tuning and alignment of LLMs. For example, Yu et al. (2021) applied DP to fine-tuning by introducing privacy guarantees for smaller, appended parameters such as LoRA and adapters. Singh et al. (2024) introduced a two-stage fine-tuning process, and Yu et al. (2024) addressed the privacy-preserving alignment challenge by ensuring DP protection for users’ prompts against labelers during the generation of preference datasets for alignment. Additionally, Wu et al. (2023) proposed applying DP to RLHF by splitting the dataset into three disjoint sets to ensure DP at each stage of RLHF. Moreover, Feng et al. (2024) investigated the vulnerability of LLMs aligned using human preference datasets to membership inference attacks (MIAs), and provided empirical evidence that DPO models are more vulnerable to MIA compared to RLHF based models.
@@
Private Alignment: To address privacy threats, the idea of Differential Privacy (DP) (Dwork et al., 2014) has been integrated into the LLM training process, including pre-training, fine-tuning, and alignment, to protect the privacy of preference data. To achieve DP, current state-of-the-art approaches utilize DP-SGD (Abadi et al., 2016), which privatizes gradients to ensure the privacy of the entire preference data throughout the training process. Previous literature has explored key aspects relevant to this study: (Chowdhury et al., 2024b) examined privacy preserving reward estimation for RLHF methods, while (Chowdhury et al., 2024a) investigated the robustness of DPO with noisy preference datasets, providing the foundational basis for this study. However, to the best of our knowledge, the concept of Label Differential Privacy (Label-DP) (Ghazi et al., 2021b) has not been explored in the context of alignment. Label-DP could enable models to achieve high utility while reducing privacy leakage, particularly when protecting the labels generated by human annotators. While DP-SGD offers privacy guarantees for the entire dataset, including the prompt $x$, LLM-generated responses $(y_{1},y_{2})$, and human-generated label $\ell^{*}$; we notice that it is primarily the labels in the preference data that pose privacy risks concerning the human labelers. Based on this critical observation, we propose a novel framework for progressively private LLM alignment that ensures a certain level of privacy for human labelers. Our framework’s primary goal is to protect human-generated labels in preference datasets while achieving a better privacy-utility trade-off. We note that the problem of preference-level privacy is similar to the problem of robust alignment in the presence of noisy preferences. Specifically, recent works, including Mitchell (2023), Chowdhury et al. (2024a) and Chowdhury et al. (2024b) study the robustness of alignment when the human-annotated labels are intrinsically noisy. The distinction, however, is the following: in our setting, the injected noise (and more importantly, the parameters of the privacy-preserving mechanisms (detailed in the next Section)) are known and can be controlled as a function of the privacy parameters.
@@ 2 Preliminaries on Alignment & Privacy
$\mathbb{E}[\textsf{loss}(x,y_{1},y_{2},\ell^{*})]$ $=\mathbb{E}_{(x,y_{1},y_{2},\ell^{*})\sim\mathcal{D}}\big{\{}\mathds{1}(\ell^{*}=1)\cdot\textsf{loss}(x,y_{1}\succ y_{2})$
$+\mathds{1}(\ell^{*}=0)\cdot\textsf{loss}(x,y_{2}\succ y_{1})\big{\}}.$ (2)
@@ 2.1 Privacy for Alignment
The notion of differential privacy (DP) Dwork et al. (2014) has been adopted in the alignment frameworks to ensure that the presence or absence of a single sample in a preference dataset does not significantly alter the outcome of the model.
@@
From Preference-level DP to Labeler-level DP: Preference-privacy protects individual labeling actions, such as rating a single prompt-response pair. However, when a labeler annotates multiple prompt-response pairs across the dataset, labeler-privacy guarantees become essential. This distinction between preference-privacy and labeler-privacy has been well recognized in the literature (McMahan et al., 2017; Liu et al., 2020; Levy et al., 2021). To extend preference-privacy guarantees to the labeler-privacy, privacy accounting and composition techniques can be adopted. For instance, if a labeler contributes to $k$ labeled examples in the dataset $\mathcal{D}$ with $(\epsilon,0)$-preference privacy, the Basic Composition theorem (Dwork et al., 2014) implies a labeler-privacy guarantee of $(k\epsilon,0)$. To limit cumulative privacy loss, it is important to operate in high-privacy regimes (i.e., with small privacy budgets). Notably, we observe that PROPS performs significantly better in such regimes, where the resulting labeler-privacy guarantees are stronger. Composition is the key technique that we use to obtain labeler-level priacy from preference-level privacy when individual labeler labels more than a single prompt-response pair. For instance, in dataset $\mathcal{D}$, if any labeler labels $k$ prompt-response pairs, with Basic Composition, $(\epsilon,0)$-Preference DP will satisfy $(k\epsilon,0)$-Labeler DP. However, more sophisticated methods, such as Advanced Composition (Dwork et al., 2014), Adaptive Composition Rogers et al. (2016), and the Moments Accountant Abadi et al. (2016) can be used to obtain tighter bounds depending on the application’s privacy requirements. With a small failure probability of $\delta^{\prime}$, Advanced composition provides ($\epsilon_{\text{Labeler}},\delta_{\text{Labeler}}$) Labeler DP where $\epsilon_{\text{Labeler}}=k_{i}\epsilon^{2}+\epsilon\sqrt{2k_{i}\log(\frac{1}{\delta^{\prime}})},$ and $\delta_{\text{Labeler}}=\delta^{\prime}$.
@@ 2.2 Limitations of Existing Approaches for Differentially Private Alignment
Though RR is simple to implement and ensures strong privacy guarantees, it introduces significant noise to the labels, which can degrade alignment quality, especially in small datasets or high-privacy regimes. Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016) ensures $(\epsilon,\delta)$-DP by gradient perturbation during alignment. In each round, the gradients $(\mathbf{\bar{g}})$ are clipped with a clipping threshold $(C)$ and perturbed with Gaussian Noise $(\mathcal{N}(0,{\sigma}^{2}{C}^{2}I))$ where, noise scale $\sigma=\sqrt{2\log(1.25/\delta)}/\epsilon$. In scenarios requiring stringent privacy constraints such as alignment, DP-SGD often reduces model utility, since it perturbs the gradient privatizing the prompt-response pairs as well as the human annotated labels. This tradeoff limits its effectiveness of alignment. These limitations of aforementioned methods highlight the need for more sophisticated approaches that achieve a better balance between privacy preservation and model performance. To address this, we propose Progressively Private Self-Alignment (PROPS), a novel framework that leverages intermediate alignment stages to improve utility while ensuring preference-level privacy. For more related works on LLM alignment and Privacy, we refer the readers to the Section A.1.
@@ 3 PROPS: Progressively Private Self-Alignment
In this Section, we present Progressively Private Self-Alignment algorithm (PROPS), which is the main technical contribution of this paper. Our key idea is to preserve the privacy of the labeler generated preferences during the multi-stage alignment process for better privacy-utility trade-off. To facilitate understanding, we first describe PROPS in a two-stage $(K=2)$ setting and the generalization to arbitrary number of stages is straightforward. We begin with the preference dataset $\mathcal{D}$ consists of $n$ samples. Each sample is represented as $(x,y_{1},y_{2},\ell^{*})$, where $x$ is the prompt, $(y_{1},y_{2})$ are the large language model (LLM) generated responses, and $\ell^{*}$ is the human labeler’s preference. We partition this dataset into two halves, denoted as $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Let us perturb the labels of each entry using the RR mechanism, i.e., the labels are flipped with probability $\gamma_{\epsilon}=1/(1+e^{\epsilon})$.
@@
Let us denote the corresponding label obtained from the model $M_{1}$ for a prompt as $\ell_{M_{1}}$. To summarize, at this point, for each tuple $(x,y_{1},y_{2})$, we have access to $(\ell_{RR},\ell_{M_{1}})$. Viewing these as two noisy observations of the ground truth $\ell^{*}$, it is natural to ask the following: can one combine these predictions to obtain a better prediction about the ground truth preference?
Note that we know the error rate of RR ($\gamma_{\epsilon}$), however, we don’t know the error rate of the model $M_{1}$ (say $\gamma_{M_{1}}$, which denotes the probability with which the model $M_{1}$ makes errors). However, if we knew the error rate of the model $M_{1}$ (or an estimate for $\gamma_{M_{1}}$), then we could then use a combining approach (e.g., the maximum-likelihood estimator or MLE) to design a potentially better estimate of the ground truth label for alignment. In fact, it is not too difficult to work out the MLE combiner using two noisy observations. Assuming that the RR noise and the noise induced by the model $M_{1}$ are independent, the MLE statistic (log-likelihood ratio) can be written and simplified as follows:
$=(-1)^{\ell_{RR}}\cdot\log\left(\frac{1-\gamma_{\epsilon}}{\gamma_{\epsilon}}\right)+(-1)^{\ell_{M_{1}}}\cdot\log\left(\frac{1-\gamma_{M_{1}}}{\gamma_{M_{1}}}\right).$ (4)
@@
The detailed proof of (8) and the fact that above estimator is unbiased is presented in Section A.3. In addition, in Section A.3, we provide experimental evidence to verify the efficacy of our proposed approach and the assumption about $M_{1}$, by comparing our proposed estimation procedure with the scenario where an oracle provides access to the true labels. Our results indicate that our procedure is valid, as the gap between the estimated and “ground-truth” $\gamma_{M}$ is consistently small across various privacy budgets.

### v2

@@ 1 Introduction
The process of aligning LLMs relies on datasets comprising prompts, LLM-generated responses, and preference labels that indicate which response aligns better with human values, collectively referred to as preference data. The alignment approaches, including reinforcement learning with human feedback (RLHF) (stiennon2020learning) and direct preference optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage), leverage these preference datasets with ranked labels provided by human annotators.
@@
Motivation for Human Preference Privacy: While preference data can significantly improve the alignment of large models with expert reasoning, it often carries deep privacy risks—particularly in domains where human feedback encodes sensitive strategies, values, or professional heuristics. This tension is most evident in several high-stakes application scenarios as we discuss next. In clinical decision support systems, for example, physicians’ feedback reflects diagnostic reasoning and treatment heuristics tied to protected health information and institutional IP; disclosure can undermine both patient privacy and clinical competitiveness (reddy2023evaluating; han2024towards). In legal and judicial settings, alignment with lawyers’ or judges’ feedback risks revealing privileged deliberations, litigation strategies, or interpretive biases that must remain confidential to preserve due process. These scenarios illustrate a central point: preference privacy is not merely a theoretical consideration but a critical requirement for deploying aligned models in the most sensitive and societally consequential domains. The privacy–utility trade-offs in such settings demand specialized alignment mechanisms that protect individual and institutional preference signals while preserving their utility for model improvement.
In wang2024gpt, a large number of consultation records were used to construct a dataset to fine-tune an LLM to act as a chatbot physician, with GPT-4 used to retain records that highlighted professionalism, explainability, and emotional support. As the model was shown to be effective, this indicates that certain preferences of a doctor’s decision-making can be identified which could lead to potential exposure of a doctor’s preferences. Also, in policy analysis (rao2023ethical), publicly available survey questions or proposals can be used to elicit LLM-generated analyses, where policymakers’ feedback reveals sensitive interpretative insights that may require protection. bakker2021reconsidering, for example, observed that placing more emphasis on politics in their surveys to participants “resulted in self-reports of personality traits that were in some cases more aligned with preexisting political preferences."
@@
Main Contributions: Motivated by the above observations, we study the problem of aligning LLMs with preference privacy. Specifically, we investigate two notions of privacy: a) preference-level privacy, and b) labeler-level privacy. Preference level privacy ensures that for any tuple $(x,y_{1},y_{2})$, where $x$ denotes the prompt and $y_{1},y_{2}$ denote the LLM-generated responses, the individual human-preference $\ell^{*}$ (which denotes whether $y_{1}$ or $y_{2}$ is preferable) does not significantly impact the aligned model. Formally, we leverage the existing notion of Differential Privacy (DP) dwork2014algorithmic, and use it to formalize the notion of $(\epsilon,\delta)$-preference-level DP, where $(\epsilon,\delta)$ represent the privacy budgets. The notion of labeler-level privacy (also commonly referred to as user-level privacy in the DP literature) hides the presence/absence of any individual human labeler and protects all the labels annotated by the labeler. We summarize and highlight the main contributions and novel aspects of this work:
  • PROPS for Private Alignment. We introduce Progressively Private Self-Alignment (PROPS), a multi-stage algorithm for alignment that improves privacy and utility. Instead of processing the entire perturbed dataset at once, PROPS divides alignment into stages. In the $k^{\text{th}}$ stage, for $(k>1)$, the model from the previous stage ($M_{k-1}$) is used along with non-private prompts, responses from a new batch, and noisy labels perturbed using Randomized Response (RR) $\ell_{\text{RR}}$. $M_{k-1}$ generates its own rankings $\ell_{M_{k-1}}$ which, combined with $\ell_{\text{RR}}$, are used to compute maximum-likelihood estimates (MLEs) for alignment. This staged process leverages intermediate models to improve preference labeling and reduce reliance on noisy labels, enhancing alignment quality while maintaining privacy. PROPS effectively balances privacy preservation with performance, offering a novel alignment framework.
  • Theoretical Insights. We study the utility-privacy tradeoffs for PROPS by analyzing the Sub-optimality gap chowdhury2024provably defined as the difference between between the weights of a non-privately trained model with the weights of a privately trained model with a privacy budget $\epsilon$ in Section 3.1 (Theorem 1). The upper bound on this gap shows that PROPS (which uses MLE combining in the second stage) is no worse than vanilla RR and performs better as long as the intermediate model gets better at predicting preferences.
@@
Consistent with standard approaches for alignment, this work focuses on the common setting of binary preferences (pairwise comparisons). The core ideas could potentially be extended to multiple preferences using techniques in Zhu2023Principled, we leave this generalization as future work.
@@
LLM Alignment: Training LLMs typically involve three key stages: pre-training, supervised fine-tuning, and alignment. Among these, alignment is particularly important as it guides LLMs to produce responses that align with societal norms and human preferences. The alignment process relies on a dataset $\mathcal{D}$ consisting of $n$ samples, each containing a prompt $x$, LLM-generated responses $(y_{1},y_{2})$, and a human-preferred label $\ell^{*}$, collectively referred to as preference data. Two conventional methods for alignment, Reinforcement Learning with Human Feedback (RLHF) (stiennon2020learning) and Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage), utilize preference datasets with ranked labels provided by human annotators. While these aligned models improve the quality of generated responses, they introduce privacy risks, particularly concerning the identity of the annotators and their associated preferences. Recent work shows membership inference attack on preference data for LLM alignment (feng2024exposing), and highlights the vulnerabilities of using human annotated preference data during alignment. To mitigate the privacy risks associated with human-annotated preference data, the notion of differential privacy (DP) (dwork2014algorithmic) has recently been explored for fine-tuning and alignment of LLMs. For example, yu2021differentially applied DP to fine-tuning by introducing privacy guarantees for smaller, appended parameters such as LoRA and adapters. singh2024whispered introduced a two-stage fine-tuning process, and yu2024privacy addressed the privacy-preserving alignment challenge by ensuring DP protection for users’ prompts against labelers during the generation of preference datasets for alignment. Additionally, wu2023privately proposed applying DP to RLHF by splitting the dataset into three disjoint sets to ensure DP at each stage of RLHF. Moreover, feng2024exposing investigated the vulnerability of LLMs aligned using human preference datasets to membership inference attacks (MIAs), and provided empirical evidence that DPO models are more vulnerable to MIA compared to RLHF based models.
@@
Private Alignment: To address privacy threats, the idea of Differential Privacy (DP) (dwork2014algorithmic) has been integrated into the LLM training process, including pre-training, fine-tuning, and alignment, to protect the privacy of preference data. To achieve DP, current state-of-the-art approaches utilize DP-SGD (abadi2016deep), which privatizes gradients to ensure the privacy of the entire preference data throughout the training process. Previous literature has explored key aspects relevant to this study: (chowdhury2024differentially) examined privacy preserving reward estimation for RLHF methods, while (chowdhury2024provably) investigated the robustness of DPO with noisy preference datasets, providing the foundational basis for this study. However, to the best of our knowledge, the concept of Label Differential Privacy (Label-DP) (ghazi2021deep) has not been explored in the context of alignment. Label-DP could enable models to achieve high utility while reducing privacy leakage, particularly when protecting the labels generated by human annotators. While DP-SGD offers privacy guarantees for the entire dataset, including the prompt $x$, LLM-generated responses $(y_{1},y_{2})$, and human-generated label $\ell^{*}$; we notice that it is primarily the labels in the preference data that pose privacy risks concerning the human labelers. Based on this critical observation, we propose a novel framework for progressively private LLM alignment that ensures a certain level of privacy for human labelers. Our framework’s primary goal is to protect human-generated labels in preference datasets while achieving a better privacy-utility trade-off. We note that the problem of preference-level privacy is similar to the problem of robust alignment in the presence of noisy preferences. Specifically, recent works, including cdpo, chowdhury2024provably and chowdhury2024differentially study the robustness of alignment when the human-annotated labels are intrinsically noisy. The distinction, however, is the following: in our setting, the injected noise (and more importantly, the parameters of the privacy-preserving mechanisms (detailed in the next Section)) are known and can be controlled as a function of the privacy parameters. Our framework can be broadly positioned within the literature on self-training, where a model ($M_{1}$) generates labels for a subsequent model ($M_{2}$). However, our core novelty is distinct. Standard self-training does not operate in a privacy-preserving context. Our contribution lies in the mechanism for combining a known privacy mechanism (RR) with the unknown-quality predictions of the $M_{1}$ model. The novelty is our method for privately estimating the intermediate model’s error rate ($\hat{\gamma}_{M_{1}}$) using the disagreement between the two noisy signals, which allows us to provably create a higher-quality private label set than using either signal alone.
@@ 2 Preliminaries on Alignment & Privacy
$\mathbb{E}[\textsf{loss}(x,y_{1},y_{2},\ell^{*})]=\mathbb{E}_{(x,y_{1},y_{2},\ell^{*})\sim\mathcal{D}}\big\{\mathds{1}(\ell^{*}=1)\cdot\textsf{loss}(x,y_{1}\succ y_{2})+\mathds{1}(\ell^{*}=0)\cdot\textsf{loss}(x,y_{2}\succ y_{1})\big\}.$ (2)
@@ 2.1 Privacy for Alignment
The notion of differential privacy (DP) dwork2014algorithmic has been adopted in the alignment frameworks to ensure that the presence or absence of a single sample in a preference dataset does not significantly alter the outcome of the model.
@@
From Preference-level DP to Labeler-level DP: Preference-privacy protects individual labeling actions, such as rating a single prompt-response pair. However, when a labeler annotates multiple prompt-response pairs across the dataset, labeler-privacy guarantees become essential. This distinction between preference-privacy and labeler-privacy has been well recognized in the literature (mcmahan2017communication; liu2020learning; levy2021learning). To extend preference-privacy guarantees to the labeler-privacy, privacy accounting and composition techniques can be adopted. For instance, if a labeler contributes to $k$ labeled examples in the dataset $\mathcal{D}$ with $(\epsilon,0)$-preference privacy, the Basic Composition theorem (dwork2014algorithmic) implies a labeler-privacy guarantee of $(k\epsilon,0)$. To limit cumulative privacy loss, it is important to operate in high-privacy regimes (i.e., with small privacy budgets). Notably, we observe that PROPS performs significantly better in such regimes, where the resulting labeler-privacy guarantees are stronger. Composition is the key technique that we use to obtain labeler-level priacy from preference-level privacy when individual labeler labels more than a single prompt-response pair. For instance, in dataset $\mathcal{D}$, if any labeler labels $k$ prompt-response pairs, with Basic Composition, $(\epsilon,0)$-Preference DP will satisfy $(k\epsilon,0)$-Labeler DP. However, more sophisticated methods, such as Advanced Composition (dwork2014algorithmic), Adaptive Composition rogers_privacy_2016, and the Moments Accountant abadi2016deep can be used to obtain tighter bounds depending on the application’s privacy requirements. With a small failure probability of $\delta^{\prime}$, Advanced composition provides ($\epsilon_{\text{Labeler}},\delta_{\text{Labeler}}$) Labeler DP where $\epsilon_{\text{Labeler}}=k_{i}\epsilon^{2}+\epsilon\sqrt{2k_{i}\log(\frac{1}{\delta^{\prime}})},$ and $\delta_{\text{Labeler}}=\delta^{\prime}$.
@@ 2.2 Limitations of Existing Approaches for Differentially Private Alignment
Though RR is simple to implement and ensures strong privacy guarantees, it introduces significant noise to the labels, which can degrade alignment quality, especially in small datasets or high-privacy regimes. Differentially Private Stochastic Gradient Descent (DP-SGD) (abadi2016deep) ensures $(\epsilon,\delta)$-DP by gradient perturbation during alignment. In each round, the gradients $(\mathbf{\bar{g}})$ are clipped with a clipping threshold $(C)$ and perturbed with Gaussian Noise $(\mathcal{N}(0,{\sigma}^{2}{C}^{2}I))$ where, noise scale $\sigma=\sqrt{2\log(1.25/\delta)}/\epsilon$. In scenarios requiring stringent privacy constraints such as alignment, DP-SGD often reduces model utility, since it perturbs the gradient privatizing the prompt-response pairs as well as the human annotated labels. This tradeoff limits its effectiveness of alignment. These limitations of aforementioned methods highlight the need for more sophisticated approaches that achieve a better balance between privacy preservation and model performance. To address this, we propose Progressively Private Self-Alignment (PROPS), a novel framework that leverages intermediate alignment stages to improve utility while ensuring preference-level privacy. For more related works on LLM alignment and Privacy, we refer the readers to the Section A.1.
@@ 3 PROPS: Progressively Private Self-Alignment
In this Section, we present Progressively Private Self-Alignment algorithm (PROPS), which is the main technical contribution of this paper. Our key idea is to preserve the privacy of the labeler generated preferences during the multi-stage alignment process for better privacy-utility trade-off. To facilitate understanding, we first describe PROPS in a two-stage $(K=2)$ setting. We begin with the preference dataset $\mathcal{D}$ consists of $n$ samples. Each sample is represented as $(x,y_{1},y_{2},\ell^{*})$, where $x$ is the prompt, $(y_{1},y_{2})$ are the large language model (LLM) generated responses, and $\ell^{*}$ is the human labeler’s preference. We partition this dataset into two halves, denoted as $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Let us perturb the labels of each entry using the RR mechanism, i.e., the labels are flipped with probability $\gamma_{\epsilon}=1/(1+e^{\epsilon})$.
@@
Let us denote the corresponding label obtained from the model $M_{1}$ for a prompt as $\ell_{M_{1}}$. To summarize, at this point, for each tuple $(x,y_{1},y_{2})$ in the partition $\mathcal{D}_{2}$, we have access to two noisy observations: its RR-perturbed label $\ell_{RR}$ (also from $\mathcal{D}_{2}$) and the model’s prediction $\ell_{M_{1}}$.
Note that we know the error rate of RR ($\gamma_{\epsilon}$), however, we don’t know the error rate of the model $M_{1}$ (say $\gamma_{M_{1}}$, which denotes the probability with which the model $M_{1}$ makes errors). However, if we knew the error rate of the model $M_{1}$ (or an estimate for $\gamma_{M_{1}}$), then we could then use a combining approach (e.g., the maximum-likelihood estimator or MLE) to design a potentially better estimate of the ground truth label for alignment. In fact, it is not too difficult to work out the MLE combiner using two noisy observations. Note that these two noisy observations are fundamentally different. The $l_{RR}$ (Privacy Noise) is from the Randomized Response (RR) mechanism, and its flip probability, $\gamma_{\epsilon}$, is known, analytic, and fixed by the privacy budget $\epsilon$. In contrast, $l_{M_{1}}$ (Model generated label) is the prediction from the Stage-1 model; its error rate, $\gamma_{M_{1}}$, is unknown, empirical, and must be estimated, as it reflects the quality of the Stage-1 alignment. Assuming that the RR noise and the noise induced by the model $M_{1}$ are independent, the MLE statistic (log-likelihood ratio) can be written and simplified as follows:
$\Lambda(\ell_{RR},\ell_{M_{1}})=\log\left(\frac{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=0)}{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=1)}\right)=(-1)^{\ell_{RR}}\cdot\log\left(\frac{1-\gamma_{\epsilon}}{\gamma_{\epsilon}}\right)+(-1)^{\ell_{M_{1}}}\cdot\log\left(\frac{1-\gamma_{M_{1}}}{\gamma_{M_{1}}}\right).$ (4)
@@
The detailed proof of (8) and the fact that above estimator is unbiased is presented in Section A.3. In addition, in Section A.3, we provide experimental evidence to verify the efficacy of our proposed approach and the assumption about $M_{1}$, by comparing our proposed estimation procedure with the scenario where an oracle provides access to the true labels. Our results indicate that our procedure is valid, as the gap between the estimated and “ground-truth" $\gamma_{M}$ is consistently small across various privacy budgets.